# Yicong Qiu    Qiye Zheng

### v1

@@ Abstract
Coarse-graining a chaotic bistable oscillator into a binary symbol sequence is a standard and effective reduction, but it often obscures the geometry of the reduced state space and the structural constraints imposed by physically meaningful stochastic evolution. We develop a self-contained two-state framework that embeds the coarse-grained left/right statistics of the driven Duffing oscillator into a $2\times 2$ density-matrix representation and models inter-well switching by a two-rate Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) generator. For diagonal states the GKSL dynamics reduces exactly to the classical two-state master equation. The density-matrix language permits an operational “Bloch half-disk” embedding of phase-space data (with coarse-graining overlap parameter $c(\varepsilon)$ quantifying partition fuzziness); the GKSL model is then fitted to the diagonal marginals (populations) of this embedded state, treating $c(\varepsilon)$ as a diagnostic rather than a dynamical variable. Within this model we derive closed-form solutions for populations and off-diagonal terms, an explicit Kraus representation (generalized amplitude damping composed with dephasing and phase rotation), and practical diagnostics that falsify or support the time-homogeneous first-order Markov assumption from symbol data (order tests, Chapman–Kolmogorov consistency, run-length statistics, and windowed stationarity checks). When diagnostics reveal higher-order memory (e.g., second-order Markov effects), we show how to extend the framework to enlarged state spaces through augmented Markov models, constructing CPTP maps via discrete-time Kraus representations; we note that continuous-time GKSL generators may not exist for all empirical transition matrices. We also provide a detailed numerical-experiment and data-analysis pipeline (with figure/table templates) for validating the framework on Duffing simulations; concrete numerical results are left for a companion computational study. The density-matrix formalism should be viewed as an organizational and algebraic convenience rather than a computational speedup or a claim of quantum-classical equivalence.
@@ 1.2 Contributions and paper organization
The paper is organized as follows. Section 2 describes the driven Duffing oscillator, Poincaré sampling, binary symbol generation, soft-partition membership functions, and the two-rate GKSL model (our methodological setup). Section 3 presents analytical results: evolution equations, closed-form solutions, steady-state characterization, discrete-time formulation, and Kraus representation with equivalence proof. Section 4 develops the diagnostic framework: order tests, Chapman–Kolmogorov consistency, run-length analysis, windowed stationarity, and model-extension protocols. Section 5 outlines numerical methods and computational validation procedures; detailed figure/table templates are provided in Appendices. Section 6 discusses scope, limitations, interpretive caveats (the two-matrix distinction $\rho_{\mathrm{emb}}$ vs. $\rho_{\mathrm{mod}}$, embedding problem for higher-order extensions, relationship to Koopman methods), and future directions. Section 7 concludes. Appendices contain derivation details, the complete Kraus equivalence proof, discussion of why pure Hamiltonian evolution cannot model switching, symmetry considerations, a summary of key formulas, and a comprehensive notation reference (Appendix F).
@@ 2.5 Soft partitions and an operational overlap parameter
The overlap parameter $c(t)$ characterizes the coarse-graining resolution but is not a dynamical observable of the reduced model. The GKSL model (Section 2.7) governs the evolution of populations $p_{L}(t),p_{R}(t)$ from symbol observations; the off-diagonal entries $\rho_{LR}$ in the density-matrix representation encode initial-state uncertainty or measurement resolution, but cannot be inferred from symbol data alone. In practice, $c(\varepsilon)$ serves as a diagnostic for partition quality: smaller $c$ indicates sharper boundaries and better approximation of the classical limit (diagonal states). The GKSL steady state $\rho^{*}$ is diagonal (Theorem 3.4), which is consistent because symbols provide no boundary-proximity information.
@@ 3.1.1 Bloch-vector form
$\dot{m}_{z}$ $=-\Gamma(m_{z}-m_{z}^{*}),\qquad m_{z}^{*}=\frac{k_{RL}-k_{LR}}{\Gamma}.$ (31)
@@ 3.2.1 Steady state
$\rho_{\mathrm{mod}}^{*}=\begin{pmatrix}p_{L}^{*}&0\\ 0&p_{R}^{*}\end{pmatrix},\qquad p_{L}^{*}=\frac{k_{RL}}{\Gamma},\quad p_{R}^{*}=\frac{k_{LR}}{\Gamma}.$ (32)
@@ 3.4 Off-diagonal decay
This defines an effective decay time $\tau_{\mathrm{off}}=1/(\Gamma/2+2\kappa)$ for off-diagonal components. The off-diagonal decay describes how initial preparation uncertainty or mixed states evolve under the reduced GKSL dynamics. This is distinct from the coarse-graining overlap $c(\varepsilon)$, which characterizes the partition geometry at a given resolution $\varepsilon$ (Section 2.3). At steady state, $\rho_{LR}^{*}=0$ because symbol data provide no boundary information; the nonzero $c(\varepsilon)>0$ for soft partitions reflects the embedding choice, not a dynamical prediction.
@@ 3.5 Discrete-time Poincaré map and transition matrix
$p_{L}^{(n+1)}=p_{L}^{*}+\bigl(p_{L}^{(n)}-p_{L}^{*}\bigr)e^{-\Gamma T}.$ (35)
@@
$P_{LR}=p_{R}^{*}(1-e^{-\Gamma T}),\qquad P_{RL}=p_{L}^{*}(1-e^{-\Gamma T}).$ (37)
@@ 4.7 Kraus representation and equivalence to generalized amplitude damping
For the two-rate model in Definition 2.17, one can construct a Kraus family by composing: (i) a generalized amplitude-damping channel encoding population relaxation toward $\rho^{*}$, (ii) a phase-damping channel encoding dephasing with rate $\kappa$, and (iii) a unitary phase rotation generated by $H$.
@@ 4.8 Kraus operators
$\Gamma=k_{LR}+k_{RL},\qquad\lambda(t)=1-e^{-\Gamma t},\qquad\eta(t)=e^{-2\kappa t},\qquad p_{R}^{*}=\frac{k_{LR}}{\Gamma},\ p_{L}^{*}=\frac{k_{RL}}{\Gamma}.$
@@
$E_{0}(t)$ $=\sqrt{p_{L}^{*}}\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda(t)}\end{pmatrix}$, $E_{1}(t)$ $=\sqrt{p_{L}^{*}}\begin{pmatrix}0&\sqrt{\lambda(t)}\\ 0&0\end{pmatrix}$, (42)
$E_{2}(t)$ $=\sqrt{p_{R}^{*}}\begin{pmatrix}\sqrt{1-\lambda(t)}&0\\ 0&1\end{pmatrix}$, $E_{3}(t)$ $=\sqrt{p_{R}^{*}}\begin{pmatrix}0&0\\ \sqrt{\lambda(t)}&0\end{pmatrix}$, (43)
@@ 5.4 Rate estimation and model validation
If Markov diagnostics pass, estimate switching rates via the inversion formulas (38) and (39). Compare model-predicted populations $p_{L}(t)$ (from solving (26)–(27) with estimated rates) against empirical populations from windowed time averages. Quantify agreement using mean absolute error or Kullback–Leibler divergence. Report estimated rates $\hat{k}_{LR}$, $\hat{k}_{RL}$, steady-state populations $p_{L}^{*}$, and relaxation time $\tau=1/\hat{\Gamma}$ with confidence intervals obtained via block bootstrap resampling (Table 2).
@@ 6.3 Model checks (Markov diagnostics)
6.4 Figure and table templates
The following placeholders are designed to be replaced by your computed outputs.
@@ 6.4 Figure and table templates
Table 3: Transition counts and one-step probabilities at sampling interval $T$. Replace the angle-bracket placeholders $\langle\cdot\rangle$ with empirical values.
@@
from $L$ $\langle N_{LL}\rangle$ $\langle N_{LR}\rangle$
from $R$ $\langle N_{RL}\rangle$ $\langle N_{RR}\rangle$
$\hat{P}_{LR}$ $\langle\hat{P}_{LR}\rangle$
$\hat{P}_{RL}$ $\langle\hat{P}_{RL}\rangle$
@@
$\hat{\Gamma}$ $\langle\hat{\Gamma}\ \pm\ \mathrm{SE}\rangle$
$\hat{k}_{LR}$ $\langle\hat{k}_{LR}\ \pm\ \mathrm{SE}\rangle$
$\hat{k}_{RL}$ $\langle\hat{k}_{RL}\ \pm\ \mathrm{SE}\rangle$
Figure 3: Population relaxation: comparison of the theoretical curve $p_{L}(t)=p_{L}^{*}+(p_{0}-p_{L}^{*})e^{-\Gamma t}$ (Theorem 3.5) to an ensemble-averaged estimate of $p_{L}(t)$ from simulation, started from biased initial conditions.
@@
Figure 6: Overlap parameter $\hat{c}(\varepsilon)$ and population $p_{L}(\varepsilon)$ versus soft-partition fuzziness $\varepsilon$. (a) Overlap $\hat{c}(\varepsilon)=\langle\sqrt{w_{L}w_{R}}\rangle$ decreases toward zero as $\varepsilon\to 0^{+}$, demonstrating convergence to hard-partition limit (Proposition 2.8). (b) Left-well population $p_{L}(\varepsilon)=\langle w_{L}\rangle$ converges to the hard-partition value as $\varepsilon\to 0^{+}$ and varies with increased fuzziness. This quantifies the sensitivity of the embedded state $\rho_{\mathrm{emb}}(\varepsilon)$ to the resolution parameter $\varepsilon$ (Section 2.3). These diagnostics characterize how the embedding choice (controlled by $\varepsilon$) affects the coarse-grained state representation, but are not predicted by the GKSL dynamics, which governs only population evolution from symbol observations (Section 2.3). The GKSL steady state is diagonal ($\rho_{LR}^{*}=0$, Theorem 3.4) because symbols contain no boundary-proximity information. Parameters: $\alpha=1.0$, $\beta=1.0$, $\delta=0.15$, $\gamma_{0}=0.3$, $\omega=1.0$.
@@ 8 Conclusion
A detailed numerical pipeline and figure/table templates were provided to support a companion computational validation study.
@@ Appendix B Proof details for the Kraus equivalence
$E_{0}\rho E_{0}^{\dagger}$ $=p_{L}^{*}\begin{pmatrix}a&c\sqrt{1-\lambda}\\ c^{*}\sqrt{1-\lambda}&b(1-\lambda)\end{pmatrix}$,
$E_{1}\rho E_{1}^{\dagger}$ $=p_{L}^{*}\begin{pmatrix}\lambda b&0\\ 0&0\end{pmatrix}$,
$E_{2}\rho E_{2}^{\dagger}$ $=p_{R}^{*}\begin{pmatrix}(1-\lambda)a&c\sqrt{1-\lambda}\\ c^{*}\sqrt{1-\lambda}&b\end{pmatrix}$,
$E_{3}\rho E_{3}^{\dagger}$ $=p_{R}^{*}\begin{pmatrix}0&0\\ 0&\lambda a\end{pmatrix}$.
@@
$(\mathcal{E}_{t}(\rho))_{LL}=(1-\lambda)a+\lambda p_{L}^{*},\qquad(\mathcal{E}_{t}(\rho))_{LR}=\sqrt{1-\lambda}\,c.$
@@
$\rho_{LL}(t)=(1-\lambda(t))\,\rho_{LL}(0)+\lambda(t)p_{L}^{*}=p_{L}^{*}+(\rho_{LL}(0)-p_{L}^{*})e^{-\Gamma t},$
@@ Appendix E Summary of key formulas
Steady state $p_{L}^{*}=k_{RL}/\Gamma$, $p_{R}^{*}=k_{LR}/\Gamma$; $\rho_{\mathrm{mod}}^{*}]_{LR}=0$
Population solution $p_{L}(t)=p_{L}^{*}+(p_{0}-p_{L}^{*})e^{-\Gamma t}$
@@
Discrete-time $P_{ij}$ $P_{LR}=p_{R}^{*}(1-e^{-\Gamma T})$, $P_{RL}=p_{L}^{*}(1-e^{-\Gamma T})$
@@ F.3 Quantum-inspired two-state representation
$\rho^{*}$ Steady-state density matrix Thm. 3.4
@@
$p_{L}^{*}$ Steady-state left-well population $k_{RL}/\Gamma$ Thm. 3.4
$p_{R}^{*}$ Steady-state right-well population $k_{LR}/\Gamma$ Thm. 3.4
@@ F.4 Pauli matrices and Bloch representation
$m_{z}^{*}$ Steady-state $z$-component $(k_{RL}-k_{LR})/\Gamma$ Eq. (31)
@@ F.10 Mathematical operators and notation
$A^{*}$ Complex conjugate or steady-state value (context-dependent) Throughout

### v3

@@ Abstract
Coarse-graining a chaotic bistable oscillator into a binary symbol sequence is a standard reduction, but it often obscures the geometry of the reduced state space and structural constraints of physically meaningful stochastic evolution. We develop a two-state framework that embeds coarse-grained left/right statistics of the driven Duffing oscillator into a $2\times 2$ density-matrix representation and models inter-well switching by a two-rate Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) generator. For diagonal states the GKSL dynamics reduces to the classical two-state master equation. The density-matrix language permits an operational “Bloch half-disk” embedding with overlap parameter $c(\varepsilon)$ quantifying partition fuzziness; the GKSL model is fitted to diagonal marginals treating $c(\varepsilon)$ as diagnostic. We derive closed-form solutions, an explicit Kraus representation (generalized amplitude damping with dephasing and rotation), and practical diagnostics for the time-homogeneous first-order Markov assumption (order tests, Chapman–Kolmogorov consistency, run-length statistics, stationarity checks). When higher-order memory appears, we extend the framework via augmented Markov models, constructing CPTP maps through discrete-time Kraus representations; continuous-time GKSL generators may not exist for all empirical transition matrices. We provide a numerical pipeline for validating the framework on Duffing simulations. The density-matrix formalism is an organizational convenience rather than claiming quantum-classical equivalence.
@@ 1.2 Contributions and paper organization
The paper is organized as follows. Section 2 describes the driven Duffing oscillator, Poincaré sampling, binary symbol generation, soft-partition membership functions, and the two-rate GKSL model (our methodological setup). Section 3 presents analytical results: evolution equations, closed-form solutions, steady-state characterization, discrete-time formulation, and Kraus representation with equivalence proof. Section 4 develops the diagnostic framework: order tests, Chapman–Kolmogorov consistency, run-length analysis, windowed stationarity, and model-extension protocols. Section 5 outlines numerical methods and computational validation procedures; detailed numerical examples are provided in Appendices. Section 6 discusses scope, limitations, interpretive caveats (the two-matrix distinction $\rho_{\mathrm{emb}}$ vs. $\rho_{\mathrm{mod}}$, embedding problem for higher-order extensions, relationship to Koopman methods), and future directions. Section 7 concludes. Appendices contain derivation details, the complete Kraus equivalence proof, discussion of why pure Hamiltonian evolution cannot model switching, symmetry considerations, a summary of key formulas, and a comprehensive notation reference (Appendix F).
@@ 2.5 Soft partitions and an operational overlap parameter
The overlap parameter $c(t)$ characterizes the coarse-graining resolution but is not a dynamical observable of the reduced model. The GKSL model (Section 2.7) governs the evolution of populations $p_{L}(t),p_{R}(t)$ from symbol observations; the off-diagonal entries $\rho_{LR}$ in the density-matrix representation encode initial-state uncertainty or measurement resolution, but cannot be inferred from symbol data alone. In practice, $c(\varepsilon)$ serves as a diagnostic for partition quality: smaller $c$ indicates sharper boundaries and better approximation of the classical limit (diagonal states). The GKSL steady state $\rho^{\infty}$ is diagonal (Theorem 3.4), which is consistent because symbols provide no boundary-proximity information.
@@ 3.1.1 Bloch-vector form
$\dot{m}_{z}$ $=-\Gamma(m_{z}-m_{z}^{\infty}),\qquad m_{z}^{\infty}=\frac{k_{RL}-k_{LR}}{\Gamma}.$ (31)
@@ 3.2.1 Steady state
$\rho_{\mathrm{mod}}^{\infty}=\begin{pmatrix}p_{L}^{\infty}&0\\ 0&p_{R}^{\infty}\end{pmatrix},\qquad p_{L}^{\infty}=\frac{k_{RL}}{\Gamma},\quad p_{R}^{\infty}=\frac{k_{LR}}{\Gamma}.$ (32)
@@ 3.4 Off-diagonal decay
This defines an effective decay time $\tau_{\mathrm{off}}=1/(\Gamma/2+2\kappa)$ for off-diagonal components. The off-diagonal decay describes how initial preparation uncertainty or mixed states evolve under the reduced GKSL dynamics. This is distinct from the coarse-graining overlap $c(\varepsilon)$, which characterizes the partition geometry at a given resolution $\varepsilon$ (Section 2.3). At steady state, $\rho_{LR}^{\infty}=0$ because symbol data provide no boundary information; the nonzero $c(\varepsilon)>0$ for soft partitions reflects the embedding choice, not a dynamical prediction.
@@ 3.5 Discrete-time Poincaré map and transition matrix
$p_{L}^{(n+1)}=p_{L}^{\infty}+\bigl(p_{L}^{(n)}-p_{L}^{\infty}\bigr)e^{-\Gamma T}.$ (35)
@@
$P_{LR}=p_{R}^{\infty}(1-e^{-\Gamma T}),\qquad P_{RL}=p_{L}^{\infty}(1-e^{-\Gamma T}).$ (37)
@@ 4.7 Kraus representation and equivalence to generalized amplitude damping
For the two-rate model in Definition 2.17, one can construct a Kraus family by composing: (i) a generalized amplitude-damping channel encoding population relaxation toward $\rho^{\infty}$, (ii) a phase-damping channel encoding dephasing with rate $\kappa$, and (iii) a unitary phase rotation generated by $H$.
@@ 4.8 Kraus operators
$\Gamma=k_{LR}+k_{RL},\qquad\lambda(t)=1-e^{-\Gamma t},\qquad\eta(t)=e^{-2\kappa t},\qquad p_{R}^{\infty}=\frac{k_{LR}}{\Gamma},\ p_{L}^{\infty}=\frac{k_{RL}}{\Gamma}.$
@@
$E_{0}(t)$ $=\sqrt{p_{L}^{\infty}}\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda(t)}\end{pmatrix}$, $E_{1}(t)$ $=\sqrt{p_{L}^{\infty}}\begin{pmatrix}0&\sqrt{\lambda(t)}\\ 0&0\end{pmatrix}$, (42)
$E_{2}(t)$ $=\sqrt{p_{R}^{\infty}}\begin{pmatrix}\sqrt{1-\lambda(t)}&0\\ 0&1\end{pmatrix}$, $E_{3}(t)$ $=\sqrt{p_{R}^{\infty}}\begin{pmatrix}0&0\\ \sqrt{\lambda(t)}&0\end{pmatrix}$, (43)
@@ 5.4 Rate estimation and model validation
If Markov diagnostics pass, estimate switching rates via the inversion formulas (38) and (39). Compare model-predicted populations $p_{L}(t)$ (from solving (26)–(27) with estimated rates) against empirical populations from windowed time averages. Quantify agreement using mean absolute error or Kullback–Leibler divergence. Report estimated rates $\hat{k}_{LR}$, $\hat{k}_{RL}$, steady-state populations $p_{L}^{\infty}$, and relaxation time $\tau=1/\hat{\Gamma}$ with confidence intervals obtained via block bootstrap resampling (Table 2).
@@ 6.3 Model checks (Markov diagnostics)
6.4 Numerical results
@@
Table 3: Transition counts and one-step probabilities at sampling interval $T=2\pi/\omega$.
@@
from $L$ $75$ $10$
from $R$ $10$ 44
$\hat{P}_{LR}$ $0.1176$
$\hat{P}_{RL}$ $0.7143$
@@
$\hat{\Gamma}$ $0.2838\pm 0.1397$
$\hat{k}_{LR}$ $0.2437\pm 0.1245$
$\hat{k}_{RL}$ $0.0401\pm 0.0207$
Figure 3: Population relaxation: comparison of the theoretical curve $p_{L}(t)=p_{L}^{\infty}+(p_{0}-p_{L}^{\infty})e^{-\Gamma t}$ (Theorem 3.5) to an ensemble-averaged estimate of $p_{L}(t)$ from simulation, started from biased initial conditions.
@@
Figure 6: Overlap parameter $\hat{c}(\varepsilon)$ and population $p_{L}(\varepsilon)$ versus soft-partition fuzziness $\varepsilon$. (a) Overlap $\hat{c}(\varepsilon)=\langle\sqrt{w_{L}w_{R}}\rangle$ decreases toward zero as $\varepsilon\to 0^{+}$, demonstrating convergence to hard-partition limit (Proposition 2.8). (b) Left-well population $p_{L}(\varepsilon)=\langle w_{L}\rangle$ converges to the hard-partition value as $\varepsilon\to 0^{+}$ and varies with increased fuzziness. This quantifies the sensitivity of the embedded state $\rho_{\mathrm{emb}}(\varepsilon)$ to the resolution parameter $\varepsilon$ (Section 2.3). These diagnostics characterize how the embedding choice (controlled by $\varepsilon$) affects the coarse-grained state representation, but are not predicted by the GKSL dynamics, which governs only population evolution from symbol observations (Section 2.3). The GKSL steady state is diagonal ($\rho_{LR}^{\infty}=0$, Theorem 3.4) because symbols contain no boundary-proximity information. Parameters: $\alpha=1.0$, $\beta=1.0$, $\delta=0.15$, $\gamma_{0}=0.3$, $\omega=1.0$.
@@ 8 Conclusion
A detailed numerical pipeline was provided to support computational validation.
Appendices
@@ Appendix B Proof details for the Kraus equivalence
$E_{0}\rho E_{0}^{\dagger}$ $=p_{L}^{\infty}\begin{pmatrix}a&c\sqrt{1-\lambda}\\ c^{*}\sqrt{1-\lambda}&b(1-\lambda)\end{pmatrix}$,
$E_{1}\rho E_{1}^{\dagger}$ $=p_{L}^{\infty}\begin{pmatrix}\lambda b&0\\ 0&0\end{pmatrix}$,
$E_{2}\rho E_{2}^{\dagger}$ $=p_{R}^{\infty}\begin{pmatrix}(1-\lambda)a&c\sqrt{1-\lambda}\\ c^{*}\sqrt{1-\lambda}&b\end{pmatrix}$,
$E_{3}\rho E_{3}^{\dagger}$ $=p_{R}^{\infty}\begin{pmatrix}0&0\\ 0&\lambda a\end{pmatrix}$.
@@
$(\mathcal{E}_{t}(\rho))_{LL}=(1-\lambda)a+\lambda p_{L}^{\infty},\qquad(\mathcal{E}_{t}(\rho))_{LR}=\sqrt{1-\lambda}\,c.$
@@
$\rho_{LL}(t)=(1-\lambda(t))\,\rho_{LL}(0)+\lambda(t)p_{L}^{\infty}=p_{L}^{\infty}+(\rho_{LL}(0)-p_{L}^{\infty})e^{-\Gamma t},$
@@ Appendix E Summary of key formulas
Steady state $p_{L}^{\infty}=k_{RL}/\Gamma$, $p_{R}^{\infty}=k_{LR}/\Gamma$; $\rho_{\mathrm{mod}}^{\infty}]_{LR}=0$
Population solution $p_{L}(t)=p_{L}^{\infty}+(p_{0}-p_{L}^{\infty})e^{-\Gamma t}$
@@
Discrete-time $P_{ij}$ $P_{LR}=p_{R}^{\infty}(1-e^{-\Gamma T})$, $P_{RL}=p_{L}^{\infty}(1-e^{-\Gamma T})$
@@ F.1 Physical system and observables
$\Omega_{L}$ Time domain when trajectory is in left well, $\{t:x(t)<0\}$ App. D
$\Omega_{R}$ Time domain when trajectory is in right well, $\{t:x(t)>0\}$ App. D
@@ F.3 Quantum-inspired two-state representation
$\rho^{\infty}$ Steady-state density matrix Thm. 3.4
@@
$p_{L}^{\infty}$ Steady-state left-well population $k_{RL}/\Gamma$ Thm. 3.4
$p_{R}^{\infty}$ Steady-state right-well population $k_{LR}/\Gamma$ Thm. 3.4
@@ F.4 Pauli matrices and Bloch representation
$m_{z}^{\infty}$ Steady-state $z$-component $(k_{RL}-k_{LR})/\Gamma$ Eq. (31)
@@ F.10 Mathematical operators and notation
$A^{*}$ Complex conjugate (Hermitian adjoint for scalars) Throughout
$A^{\infty}$ Steady-state (long-time) value Throughout